# GSPR: Multimodal Place Recognition Using 3D Gaussian Splatting for Autonomous Driving

### v1

@@ Abstract
Place recognition is a crucial module to ensure autonomous vehicles obtain usable localization information in GPS-denied environments. In recent years, multimodal place recognition methods have gained increasing attention due to their ability to overcome the weaknesses of unimodal sensor systems by leveraging complementary information from different modalities. However, challenges arise from the necessity of harmonizing data across modalities and exploiting the spatio-temporal correlations between them sufficiently. In this paper, we propose a 3D Gaussian Splatting-based multimodal place recognition neural network dubbed GSPR. It explicitly combines multi-view RGB images and LiDAR point clouds into a spatio-temporally unified scene representation with the proposed Multimodal Gaussian Splatting. A network composed of 3D graph convolution and transformer is designed to extract high-level spatio-temporal features and global descriptors from the Gaussian scenes for place recognition. We evaluate our method on the nuScenes dataset, and the experimental results demonstrate that our method can effectively leverage complementary strengths of both multi-view cameras and LiDAR, achieving SOTA place recognition performance while maintaining solid generalization ability. Our open-source code is available at https://github.com/QiZS-BIT/GSPR.
Index Terms:
Place Recognition, Sensor Fusion, 3D Gaussian Splatting.
@@ I Introduction
Given an observation from sensors at the current moment (query), place recognition needs to determine which location in the global map (database) the observation corresponds to. Place recognition is an important module in most navigation systems, capable of correcting accumulated drift in SLAM algorithms and often serving as the first step in global localization. In autonomous driving systems, cameras are commonly used for vision-based place recognition (VPR), providing rich semantics and texture information [1, 2, 3]. However, vision features extracted from camera images exhibit lower stability and result in suboptimal recognition accuracy when facing variations in lighting, seasons, and weather in large-scale outdoor environments. In contrast, LiDAR sensors show high stability against these factors, leading to more robust LiDAR-based place recognition (LPR) in large-scale outdoor scenes [4, 5, 6]. However, the recognition performance of LPR is still limited by the natural sparsity of LiDAR point clouds, and the lack of texture and semantic information.
@@
In recent years, multimodal place recognition (MPR) methods such as MinkLoc++ [7] and LCPR [8] have demonstrated the potential advantages of fusing data from complementary camera and LiDAR modalities, attracting more research interests. However, most MPR methods independently extract features from primitive representations of scenes, i.e., raw images and point clouds, and implement unexplainable feature-level fusion. This results in insufficient exploitation of the spatio-temporal correlations between different modalities. Therefore, how to effectively fuse multimodal sensor data into a unified scene representation and fully extract multimodal spatio-temporal correlations remains a topic worthy of further research. Recently, 3D Gaussian Splatting (3D-GS) [9] is proposed to construct an explicit scene representation using 3D Gaussians, ensuring fast rendering of novel view while effectively capturing accurate geometry information. By aggregating temporally continuous observations from multiple views, 3D-GS constructs spatial structural representations comprehensively, providing the possibility for spatio-temporal fusion of multimodal place recognition.
@@
Figure 1: Effectively integrating different modalities is crucial for leveraging multimodal data. GSPR harmonizes multi-view RGB images and LiDAR point clouds into a unified scene representation based on Multimodal Gaussian Splatting. 3D graph convolution and transformer are utilized to extract both local and global spatio-temporal information embedded in the scene. This approach enables the generation of discriminative descriptors for place recognition.
@@
In this paper, we propose a 3D Gaussian Splatting-based multimodal place recognition method namely GSPR, as shown in Fig. 1. We first design a Multimodal Gaussian Splatting (MGS) method to represent autonomous driving scenarios. We utilize LiDAR point clouds as a prior for the initialization of Gaussians, which helps to address the failures of structure-from-motion (SfM) in such environments. In addition, a mixed masking mechanism is employed to remove unstable features less valuable for place recognition. By doing so, we fuse multimodal data into a spatio-temporally unified Gaussian scene representation. We then downsample the unordered Gaussians within each scene into a set of voxels through voxel partitioning, and develop a network based on 3D graph convolution and transformer to extract high-level spatio-temporal features for generating discriminative descriptors for place recognition. Through the proposed Multimodal Gaussian Splatting, we fuse multimodal data into a unified explicit scene representation, providing a foundation for spatio-temporal fusion of multimodal place recognition.
@@
  • We propose Multimodal Gaussian Splatting method to harmonize multi-view camera and LiDAR data into a spatio-temporally unified explicit scene representation.
  • We propose GSPR, a novel MPR network equipped with 3D graph convolution and transformer to aggregate high-level local and global spatio-temporal information inherent in the MGS scene representation.
  • Extensive experimental results demonstrate that our method outperforms the state-of-the-art unimodal and multimodal methods on place recognition performance while showing a solid generalization ability on unseen driving scenarios.
@@ II-A Scene Representation in Place Recognition
Place recognition is a classic topic in the fields of robotics and computer vision, and there have been various types of traditional methods based on handcrafted descriptors [10, 1, 11]. With the rapid development of deep learning, an increasing number of learning-based approaches [2, 12, 4, 5, 6] have been proposed and overall present better recognition performance than traditional counterparts.
In place recognition tasks, autonomous vehicles perceive the environment through cameras or LiDAR sensors and attempt to build a reasonable scene representation corresponding to the place where the vehicle is situated. The input form of place recognition methods is closely related to the type of sensors. Most vision-based place recognition methods [2, 12, 13, 14] treat RGB images as trivial scene representations. NetVLAD [2] aggregates features from RGB images into global descriptors. Building on this, SPE-NetVLAD [12] and Patch-NetVLAD [3] introduces a focus on local image features, enhancing the place recognition performance in complex environments. R2Former [13] directly inputs RGB images into transformers for image retrieval and re-ranking. SALAD [14] aggregates image features using optimal transport algorithm, producing discriminative descriptors. LiDAR or Radar-based place recognition [15, 16, 17, 18, 8], on the other hand, uses sparse point cloud representations from range detection to depict driving scenes. For instance, PointNetVLAD [4] uses submaps obtained by stacking LiDAR point clouds as the scene representation, Autoplace [16] uses BEV views constructed from multi-view radars to capture structural information of the scene, OverlapNet [19] obtains dense depth information by projecting unordered LiDAR point clouds into range images, and CVTNet [18] combines multi-layer BEVs and range images to alleviate the information loss of 3D point cloud projection. However, these primitive scene representations typically fail to adequately reflect the spatio-temporal correlation within the scene. Our method, utilizing 3D-GS, generates unified explicit scene representations, which encompass richer and more harmonized spatio-temporal information compared to primitive representations.
@@ II-B Multimodal Place Recognition
Recently, multimodal place recognition has aroused great interest due to its ability to leverage the complementary advantages of multiple sensors. PIC-Net proposed by Lu et al. [20] uses global channel attention to fuse descriptors from point clouds and images. MinkLoc++ [7] concatenates point cloud descriptors from sparse convolutions with image descriptors from pre-trained ResNet Blocks. AdaFusion proposed by Lai et al. [21] adjusts the weights of different modalities in the global descriptor through a weight generation branch. LCPR introduced by Zhou et al. [8] employs multi-scale attention to explore inner feature correlations between different modalities during feature extraction. EINet proposed by Xu et al. [22] introduces a novel multimodal fusion strategy that supervises image feature extraction with LiDAR depth maps and enhances LiDAR point clouds with image texture information. It is notable that most MPR methods extract features from each modality independently or perform inexplicable feature-level fusion. In contrast, our approach employs Multimodal Gaussian Splatting to integrate multimodal data into a unified explicit scene representation, allowing for a more thorough exploitation of the spatio-temporal correlations between different modalities.
@@
3D-GS performs well in static, bounded small scenes, but faces limitations such as scale uncertainty and training view overfitting in autonomous driving scenarios. To address these challenges, Yan et al. [23] propose Street Gaussian, which uses LiDAR point prior and introduces 4D spherical harmonics to represent dynamic objects. Zhou et al. [24] introduce Driving Gaussian, integrating an incremental static Gaussian model with a composite dynamic Gaussian graph for scene reconstruction. Following this, S3Gaussian, proposed by Huang et al. [25], attempts to eliminate the reliance on annotated data by introducing a multi-resolution hex plane for self-supervised foreground-background decomposition. DHGS proposed by Shi et al. [26] uses a signed distance field to supervise the geometric attributes of road surfaces. Inspired by these works, we propose Multimodal Gaussian Splatting, leveraging multimodal data and the proposed mixed masking mechanism, to provide stable and geometrically accurate reconstruction results of autonomous driving scenes for place recognition.
@@ III-A Multimodal Gaussian Splatting
As illustrated in Fig. 3, we introduce Multimodal Gaussian Splatting for autonomous driving scene reconstruction. The method processes multimodal data through the Image Branch and the LiDAR Branch, and then integrates different modalities into a spatio-temporally unified explicit scene representation through Gaussian Optimization. The Image Branch utilizes sequential multi-view RGB images as input, generating dynamic and static masks through the mixed masking mechanism. The LiDAR Branch complements the LiDAR coverage of distant landscapes to mitigate overfitting, and provide a LiDAR prior for the initialization of Gaussians.
III-A1 LiDAR Prior
The vanilla 3D-GS uses SfM to reconstruct point clouds for initializing the Gaussian model. However, in autonomous driving scenarios, SfM can fail due to the complexity of the scene, illumination changes, and the high-speed movement of the ego vehicle. To address this, we introduce LiDAR point clouds for initializing the position of Gaussians following [23, 24]. Using LiDAR point as position prior, the distribution of 3D Gaussian can be represented as:
@@ III-A1 LiDAR Prior
where $C^{p_{i}^{\scriptscriptstyle\text{L}}}_{\text{rgb}}$ is the corresponding color of LiDAR point $p_{i}^{\scriptscriptstyle\text{L}}$, $I$ represents the image, $K_{\text{intr}}$ and $R$ are the intrinsic parameters and extrinsic parameters of the camera corresponding to image $I$, while $F$ denotes the set of LiDAR points within the frustum of the camera.
In addition, we filter the ground points from the LiDAR point cloud and employ ground-truth annotations for object bounding box erasing, in order to ensure high-quality reconstruction of the static background.
III-A2 Overfitting Mitigation
Unlike bounded scenarios that the vanilla 3D-GS can trivially render, autonomous driving scenes present challenges due to their boundlessness and sparse distribution of training views. This scarcity of supervision signals results in overfitting of training views, leading to floating artifacts and misalignment of geometric structures.
@@ III-A2 Overfitting Mitigation
Figure 3: The pipeline of Multimodal Gaussian Splatting. Processed LiDAR point clouds supply prior information for initializing the Gaussian model. RGB image sequences are utilized to generate distinct masks that supervise the Gaussian optimization process, following semantic segmentation and mixed masking mechanism. After iterative Gaussian optimization, the multimodal data are integrated into a unified MGS scene representation.
III-A3 Mixed Masking Mechanism
In autonomous driving scenes, there are environmental features that exhibit instability over time and contain less valuable information for place recognition. Therefore, we propose the mixed masking mechanism focusing on reconstructing only the stable parts during the Gaussian optimization process.
We employ Mask2Former [28], pre-trained on the Cityscapes dataset [29], as our semantic segmentation module to generate semantic labels for the training images. By integrating semantic labels with 2D ground-truth annotations, we can obtain instance-level mask representations. In light of the nature of unstable environmental features, we categorize the masked regions into static masks (e.g., sky and road surfaces) and dynamic masks (e.g., vehicles and pedestrians), each playing distinct roles during the Gaussian optimization process. The static masks are utilized as a criterion for training image culling. Areas of the training images covered by the static masks are overlaid with the background color of the 3D-GS renderer, serving to restrict the generation of Gaussians. Conversely, the dynamic masks encompass dynamic objects within the scenarios. Notably, directly culling the shadow areas of these dynamic objects may result in unnecessary information loss. Therefore, we adopt a loss detach strategy, omitting the $\mathcal{L}_{\text{MGS}}$ loss [9] (see Eq. (10)) for the masked regions during the Gaussian optimization process. This strategy mitigates the negative effects of dynamic objects and simultaneously maintains enough supervision for large-scale reconstruction compared to directly filtering out frames with dynamic objects.
As demonstrated in Fig. 4, our proposed mixed masking mechanism effectively masks out unstable features. Additionally, the employment of LiDAR prior and the adaption of overfitting mitigation techniques contribute to maintaining a consistent scale and accurate geometric structure of the reconstructed scene. Consequently, our proposed MGS exhibits enhanced novel view synthesis capabilities compared to the vanilla 3D-GS in autonomous driving scenarios, providing MGS scene representations with spatio-temporal consistency to describe places.
@@ III-B Global Descriptor Generator
Global Descriptor Generator is used to extract distinctive place recognition descriptors from the proposed MGS representations. To extract the high-level spatio-temporal features, we first voxelize the MGS scene, and then extract local and global features through a backbone network composed of 3D graph convolutions [30] and transformer [31] module. Finally, the spatio-temporal features are fed into NetVLAD-MLPs combos [4] and aggregated into discriminative descriptors.
III-B1 Voxel Partition and Encoding
@@
To tackle the disordered distribution of Gaussians, we first organize the MGS scene into a form that facilitates feature extraction through voxelization. Denote $\mathcal{G}=\{g_{m}=[x_{m}^{\scriptscriptstyle\text{G}},y_{m}^{\scriptscriptstyle\text{G}},z_{m}^{\scriptscriptstyle\text{G}},s_{m},q_{m},sh_{m},\alpha_{m}]^{\mathrm{T}}\in\mathbb{R}^{59}\}_{m=1...M}$ as a MGS scene, where $g_{m}$ represents the $m$-th Gaussian in the scene, $x_{m}^{\scriptscriptstyle\text{G}},y_{m}^{\scriptscriptstyle\text{G}},z_{m}^{\scriptscriptstyle\text{G}}$ denote the position of the Gaussian, $s_{m}$ means the scale matrix, $q_{m}$ is the quaternion, $sh_{m}$ means the SH coefficents, and $\alpha_{m}$ denotes the opacity. Inspired by [32], we subdivide the space into voxels in cylindrical coordinates, to ensure the uniformity of the partitioning of the Gaussian model. Considering the position property of the Gaussians, we allocate them to the corresponding voxels using the voxel partition implementation of [33], converting the Gaussian model with sizes of $M\times 59$ to $N\times H\times 59$, where $N$ is the number of voxels, and $H$ is the maximum number of Gaussians within each voxel.
Let $\mathcal{V}=\{g_{h}=[x_{h}^{\scriptscriptstyle\text{G}},y_{h}^{\scriptscriptstyle\text{G}},z_{h}^{\scriptscriptstyle\text{G}},s_{h},q_{h},sh_{h},\alpha_{h}]^{\mathrm{T}}\in\mathbb{R}^{59}\}_{h=1...H}$ as a non-empty voxel containing $H$ Gaussians, we use MeanVFE [34] to encode the voxel feature to ensure the real-time performance and usability of the network:
$\overline{\mathcal{V}}=\{\frac{\sum\limits_{h=1}^{H}\mu_{h}^{\scriptscriptstyle\text{G}}}{H},\frac{\sum\limits_{h=1}^{H}s_{h}}{H},\frac{\sum\limits_{h=1}^{H}q_{h}}{H},\frac{\sum\limits_{h=1}^{H}sh_{h}}{H},\frac{\sum\limits_{h=1}^{H}\alpha_{h}}{H}\}$ (3)
@@
where $\overline{\mathcal{V}}$ denotes the encoded voxel, and $\mu^{\scriptscriptstyle\text{G}}_{h}$ means the position $(x_{h}^{\scriptscriptstyle\text{G}},y_{h}^{\scriptscriptstyle\text{G}},z_{h}^{\scriptscriptstyle\text{G}})$ of the $h$-th Gaussian in the voxel. After the voxel encoding operation, the voxel set of shape $N\times H\times 59$ is encoded into an input form of $N\times 59$. We denote the encoded MGS scene representation as $\overline{\mathcal{G}}=\{\overline{g}_{n}=[\overline{x}_{n}^{\scriptscriptstyle\text{G}},\overline{y}_{n}^{\scriptscriptstyle\text{G}},\overline{z}_{n}^{\scriptscriptstyle\text{G}},\overline{s}_{n},\overline{q}_{n},\overline{sh}_{n},\overline{\alpha}_{n}]^{\mathrm{T}}\in\mathbb{R}^{59}\}_{n=1...N}$. Ultimately, the voxel downsampling operation imparts orderliness to the Gaussian scene and reduces the number of Gaussians that need to be processed.
III-B2 3D Graph Convolution
@@
Inspired by the successful application of graph convolution in place recognition [35, 15], we use a 3D-GCN-based [30] graph convolution backbone network to fully exploit the local features in the scene.
@@
III-B3 Transformer Module
We use transformers following the previous works [36, 6] to boost place recognition performance. Compared to these works that focus primarily on self-attention mechanisms without incorporating positional embedding, we propose using a learnable positional embedding to integrate spatial correlations from the feature graph into the latent feature space of the transformer.
The architecture of the transformer module is depicted in Fig. 5. To enable the transformer to capture the spatial correlations embedded in the feature graph, we explicitly encode the coordinates of the feature graph $p_{\mathrm{feat}}^{i}$ using learnable positional embedding:
$\bm{r}_{i}=\mathrm{MLP}\left(\mathrm{ReLU}\left(\mathrm{MLP}(p_{\text{feat}}^{i})\right)\right)$ (8)
where $\bm{r}_{i}$ is the positional embedding of feature $\bm{f}(p_{\text{feat}}^{i})$.
After adding the positional embeddings to the features and performing feature fusion through 3D graph convolution layers, we feed the position-encoded features into multi-head attention to fully extract the global spatio-temporal information in the scene. The self-attention mechanism can be formulated as:
@@ III-B3 Transformer Module
TABLE I: Comparision of place recognition performance on the BS, SON and SHV split
@@
NetVLAD [2] $\times$ V 69.94 78.51 83.57 73.42 83.04 87.09 67.28 79.20 83.49
OverlapTransformer [6] $\times$ L 66.62 86.26 89.76 78.03 94.19 97.22 77.13 95.12 97.26
@@
OT+SeqNet [6, 37] $\checkmark$ L 72.61 87.50 91.29 90.38 97.47 98.73 93.88 98.47 98.78
SeqOT [17] $\checkmark$ L 78.68 89.76 93.41 94.70 98.23 98.74 96.34 98.48 99.39

### v2

@@ Abstract
Place recognition is a crucial component that enables autonomous vehicles to obtain localization results in GPS-denied environments. In recent years, multimodal place recognition methods have gained increasing attention. They overcome the weaknesses of unimodal sensor systems by leveraging complementary information from different modalities. However, most existing methods explore cross-modality correlations through feature-level or descriptor-level fusion, suffering from a lack of interpretability. Conversely, the recently proposed 3D Gaussian Splatting provides a new perspective on multimodal fusion by harmonizing different modalities into an explicit scene representation. In this paper, we propose a 3D Gaussian Splatting-based multimodal place recognition network dubbed GSPR. It explicitly combines multi-view RGB images and LiDAR point clouds into a spatio-temporally unified scene representation with the proposed Multimodal Gaussian Splatting. A network composed of 3D graph convolution and transformer is designed to extract spatio-temporal features and global descriptors from the Gaussian scenes for place recognition. Extensive evaluations on three datasets demonstrate that our method can effectively leverage complementary strengths of both multi-view cameras and LiDAR, achieving SOTA place recognition performance while maintaining solid generalization ability. Our open-source code will be released at https://github.com/QiZS-BIT/GSPR.
@@ I Introduction
Given an observation from sensors at the current moment (query), place recognition needs to determine which location in the global map (database) the observation corresponds to. Place recognition is an important module in most navigation systems, capable of correcting accumulated drift in SLAM algorithms and often serving as the first step in global localization. In autonomous driving systems, cameras are commonly used for vision-based place recognition (VPR), providing rich semantics and texture information [1, 2, 3]. However, vision features extracted from camera images exhibit lower stability and result in suboptimal recognition accuracy when facing variations in lighting, seasons, and weather in large-scale outdoor environments. In contrast, LiDAR sensors show high stability against these factors, leading to more robust LiDAR-based place recognition (LPR) [4, 5, 6]. However, the recognition performance of LPR is still limited by the natural sparsity of LiDAR point clouds, and the lack of texture and semantic information.
@@
In recent years, multimodal place recognition (MPR) methods such as MinkLoc++ [7] and LCPR [8] have demonstrated the potential advantages of fusing data from complementary camera and LiDAR modalities, attracting more research interests. Some MPR methods extract features for each modality independently, followed by descriptor-level fusion. Others generate multimodal descriptors through modality-wise feature-level fusion. However, these methods suffer from a lack of interpretability, which limits insight into cross-modal interactions. Recently, the introduction of 3D Gaussian Splatting (3D-GS) [9] provides a new perspective on multimodal fusion. It is proposed to construct an explicit scene representation using 3D Gaussians, effectively capturing the geometry information of the scene. By aggregating temporally continuous observations from multiple views, 3D-GS comprehensively constructs spatial structure representations, providing the possibility of explicable spatio-temporal fusion of multimodal place recognition.
@@
Figure 1: Effectively integrating different modalities is crucial for leveraging multimodal data. GSPR harmonizes multi-view RGB images and LiDAR point clouds into a unified scene representation based on Multimodal Gaussian Splatting. 3D graph convolution and transformer are utilized to extract both local and global spatio-temporal information embedded in the scene, ultimately generating discriminative descriptors.
@@
In this paper, we propose a 3D Gaussian Splatting-based multimodal place recognition method namely GSPR, as shown in Fig. 1. We first design a Multimodal Gaussian Splatting (MGS) method to represent autonomous driving scenarios. We utilize LiDAR point clouds as a prior for the initialization of Gaussians, which helps to address the failures of structure-from-motion (SfM) in such environments. In addition, a mixed masking mechanism is employed to remove unstable features less valuable for place recognition. By taking advantage of the attribute updates and adaptive density control strategies in the Gaussian Optimization process, we can obtain Gaussian scenes that complement the advantages of each modality. In the explicit scene representations, the Gaussians are densely and uniformly distributed, reflecting the fine-grained geometric structure of the scene. Additionally, the Gaussians encode rich semantic and texture information from the images. We then downsample the unordered Gaussians through voxel partitioning, and develop a network based on 3D graph convolution and transformer to extract high-level spatio-temporal features for generating discriminative descriptors for place recognition. Through the proposed MGS, we fuse multimodal data into a unified explicit scene representation, providing the basis for multimodal place recognition.
@@
  • We propose Multimodal Gaussian Splatting method to harmonize multi-view camera and LiDAR data into explicit scene representations suitable for place recognition.
  • We propose GSPR, a novel MPR network equipped with 3D graph convolution and transformer to aggregate local and global spatio-temporal information inherent in the MGS scene representation.
  • Extensive experimental results on three datasets demonstrate that our method outperforms the state-of-the-art unimodal and multimodal methods on place recognition performance while showing a solid generalization ability on unseen driving scenarios.
@@ II-A Scene Representation in Place Recognition
Place recognition is a classic topic in the fields of robotics and computer vision, and there have been various types of traditional methods based on handcrafted descriptors [10, 1, 11]. With the rapid development of deep learning, an increasing number of learning-based approaches [2, 4, 7, 6] have been proposed and overall present better recognition performance than traditional counterparts.
In place recognition tasks, autonomous vehicles perceive the environment through cameras or LiDAR sensors and attempt to build a reasonable scene representation corresponding to the place where the vehicle is situated. The input form of place recognition methods is closely related to the type of sensors. Most vision-based place recognition methods [2, 12, 13, 14] treat RGB images as trivial scene representations. NetVLAD [2] aggregates features from RGB images into global descriptors. To enhance robustness to appearance changes, Delta Descriptors [12] constructs change-based descriptors using sequential images. JIST [15] leverages a large uncurated set of images to mitigate the issue of limited sequential data, achieving robust place recognition. LiDAR or Radar-based place recognition [16, 17, 18, 19] represents the scene as a point cloud or its various derived forms. For instance, PointNetVLAD [4] uses submaps obtained by stacking LiDAR point clouds as the scene representation, Autoplace [17] uses BEV views constructed from multi-view radars to capture structural information of the scene, OverlapNet [20] obtains dense depth information by projecting unordered LiDAR point clouds into range images, BVMatch [21] and BEVPlace [22] achieve efficient place recognition and pose estimation through LiDAR BEVs, and CVTNet [19] combines multi-layer BEVs and range images to alleviate the information loss of 3D point cloud projection. These primitive scene representations from different modalities each have their own advantages and disadvantages. Our method, using 3D-GS, generates unified explicit scene representations that harmonize different modalities, allowing the representation of both the fine-grained geometric structure and the texture information.
@@ II-B Multimodal Place Recognition
Recently, multimodal place recognition has aroused great interest due to its ability to leverage the complementary advantages of multiple sensors. MinkLoc++ [7] concatenates point cloud descriptors from sparse convolutions with image descriptors from pre-trained ResNet Blocks. AdaFusion [23] adjusts the weights of different modalities in the global descriptor through a weight generation branch. LCPR [8] employs multi-scale attention to explore inner feature correlations between different modalities during feature extraction. EINet [24] introduces a novel multimodal fusion strategy that supervises image feature extraction with LiDAR depth maps and enhances LiDAR point clouds with image texture information. It is notable that most MPR methods fuse abstracted descriptor vectors or integrate features from different modalities to harmonize multimodal data. However, the process by which the two modalities complement and integrate remains neither explicit nor explainable. In our proposed GSPR, we instead explicitly fuse spatio-temporal information from different modalities by Multimodal Gaussian Splatting. The distributions and properties of the optimized Gaussians can reflect the rationality of harmonizing multimodal data, allowing for explicit and thorough exploitation of the spatio-temporal correlations between different modalities.
@@
3D-GS performs well in static, bounded small scenes, but faces limitations such as scale uncertainty and training view overfitting in autonomous driving scenarios. To address these challenges, Street Gaussian [25] uses LiDAR point prior and introduces 4D spherical harmonics to represent dynamic objects. Driving Gaussian [26] integrates an incremental static Gaussian model with a composite dynamic Gaussian graph for scene reconstruction. Following this, S3Gaussian [27] attempts to eliminate the reliance on annotated data by introducing a multi-resolution hex plane for self-supervised foreground-background decomposition. DHGS [28] uses a signed distance field to supervise the geometric attributes of road surfaces. Inspired by these works, we propose Multimodal Gaussian Splatting, leveraging multimodal data and the proposed mixed masking mechanism, to provide stable and geometrically accurate reconstruction results of autonomous driving scenes suitable for place recognition.
@@ III-A Multimodal Gaussian Splatting
As illustrated in Fig. 3, we introduce Multimodal Gaussian Splatting for autonomous driving scene reconstruction. The method processes multimodal data through the Image Branch and the LiDAR Branch, and then integrates different modalities into a spatio-temporally unified explicit scene representation through Gaussian Optimization. This provides a scene representation with a larger area of coverage and a more uniform distribution than the LiDAR point cloud. Additionally, each Gaussian encodes the features and texture information corresponding to the splatting region in the image, ultimately enabling explicit spatio-temporal fusion of multimodal data.
LiDAR prior. The vanilla 3D-GS uses SfM to reconstruct point clouds for initializing the Gaussian model. However, in autonomous driving scenarios, SfM can fail due to the complexity of the scene, illumination changes, and the high-speed movement of the ego vehicle. To address this, we introduce LiDAR point clouds for initializing the position of Gaussians following [25, 26]. Using LiDAR point as position prior, the distribution of 3D Gaussian can be represented as:
@@
where $C^{p_{i}^{\scriptscriptstyle\text{L}}}_{\text{rgb}}$ is the corresponding color of LiDAR point $p_{i}^{\scriptscriptstyle\text{L}}$, $I$ represents the image, $K_{\text{intr}}$ denotes the intrinsic parameters, while $R$ and $t$ represent the extrinsic parameters, associated with the camera corresponding to the image $I$, while $F$ denotes the set of LiDAR points within the frustum of the camera.
In addition, we filter the ground points from the LiDAR point cloud and employ 3D annotations for object bounding box erasing, in order to ensure high-quality reconstruction of the static background.
Overfitting mitigation. Unlike bounded scenarios that the vanilla 3D-GS can trivially render, autonomous driving scenes present challenges due to their boundlessness and sparse distribution of training views. This scarcity of supervision signals results in overfitting of training views, leading to floating artifacts and misalignment of geometric structures.
@@
Figure 3: The Multimodal Gaussian Splatting (MGS) pipeline initializes the Gaussians using processed LiDAR point clouds as prior information. RGB image sequences generate masks to guide Gaussian optimization through semantic segmentation and mixed masking. After iterative optimization, the multimodal data are integrated into a unified MGS scene representation.
Mixed masking mechanism. In autonomous driving scenes, there are environmental features that exhibit instability over time and contain less valuable information for place recognition. Therefore, we propose the mixed masking mechanism focusing on reconstructing only the stable parts during the Gaussian optimization process.
In light of the varying nature of unstable environmental features, we categorize the masked regions based on semantic information into static masks (e.g., sky and road surfaces) and dynamic masks (e.g., vehicles and pedestrians). We utilize a pre-trained Mask2Former [30] semantic segmentation network and 3D annotations to generate these two types of masks. The static masks are generated based on the semantic segmentation results. Areas of the training images covered by the static masks are overlaid with the background color of the 3D-GS renderer, serving to restrict the generation of Gaussians. The regions covered by the dynamic masks are generated through a two-step process. Firstly, 3D annotations are projected on images to obtain 2D bounding boxes. Then, pixels within the 2D bounding boxes that have the same semantic categories as the 3D annotations are selected, ensuring that static background areas are minimally masked. As directly culling the shadow areas of these dynamic objects may result in unnecessary information loss, we adopt a loss detach strategy, omitting the gradients for the masked regions during the Gaussian optimization process. This strategy mitigates the negative effects of dynamic objects and simultaneously maintains enough supervision for large-scale reconstruction.
As demonstrated in Fig. 4, our proposed mixed masking mechanism effectively masks out unstable features. Additionally, the employment of LiDAR prior and the adaption of overfitting mitigation techniques contribute to maintaining a consistent scale and accurate geometric structure of the reconstructed scene. Consequently, our MGS effectively reconstructs Gaussian scenes suitable for place recognition.
@@ III-B Global Descriptor Generator
Global Descriptor Generator is used to extract distinctive global descriptors from the proposed MGS representations. To extract the high-level spatio-temporal features, we first voxelize the MGS scene, and then extract local and global features through a backbone network composed of 3D graph convolutions [31] and transformer [32] module. Finally, the spatio-temporal features are fed into NetVLAD-MLPs combos [4] and aggregated into discriminative descriptors.
@@
Voxel partition and encoding. To tackle the disordered distribution of Gaussians, we first organize the MGS scene into a form that facilitates feature extraction through voxelization. Denote $\mathcal{G}=\{g_{m}=[x_{m}^{\scriptscriptstyle\text{G}},y_{m}^{\scriptscriptstyle\text{G}},z_{m}^{\scriptscriptstyle\text{G}},s_{m},q_{m},sh_{m},\alpha_{m}]^{\mathrm{T}}\in\mathbb{R}^{59}\}_{m=1...M}$ as a MGS scene, where $g_{m}$ represents the $m$-th Gaussian in the scene, $x_{m}^{\scriptscriptstyle\text{G}},y_{m}^{\scriptscriptstyle\text{G}},z_{m}^{\scriptscriptstyle\text{G}}$ denote the position of the Gaussian, $s_{m}$ means the scale matrix, $q_{m}$ is the quaternion, $sh_{m}$ means the SH coefficents, and $\alpha_{m}$ denotes the opacity. Inspired by [33], we subdivide the space into voxels in cylindrical coordinates, to ensure the uniformity of the partitioning of the Gaussian scene. Subsequently, we allocate the Gaussians to the corresponding voxels through voxel partitioning, converting the Gaussian model with sizes of $M\times 59$ to $N\times H\times 59$, where $N$ is the number of voxels, and $H$ is the maximum number of Gaussians within each voxel.
@@
Let $\mathcal{V}=\{g_{h}=[x_{h}^{\scriptscriptstyle\text{G}},y_{h}^{\scriptscriptstyle\text{G}},z_{h}^{\scriptscriptstyle\text{G}},s_{h},q_{h},sh_{h},\alpha_{h}]^{\mathrm{T}}\in\mathbb{R}^{59}\}_{h=1...H}$ as a non-empty voxel containing $H$ Gaussians. Inspired by [34], we encode the voxel features by computing the mean of each attribute of the Gaussians within the voxel, to ensure the real-time performance and usability of the network. After the voxel encoding operation, the voxel set of shape $N\times H\times 59$ is encoded into an input form of $N\times 59$. We denote the encoded MGS scene representation as $\overline{\mathcal{G}}=\{\overline{g}_{n}=[\overline{x}_{n}^{\scriptscriptstyle\text{G}},\overline{y}_{n}^{\scriptscriptstyle\text{G}},\overline{z}_{n}^{\scriptscriptstyle\text{G}},\overline{s}_{n},\overline{q}_{n},\overline{sh}_{n},\overline{\alpha}_{n}]^{\mathrm{T}}\in\mathbb{R}^{59}\}_{n=1...N}$. Ultimately, voxel downsampling brings order to the Gaussian scene and reduces the computational burden.
@@
3D graph convolution. Inspired by the successful application of graph convolution in place recognition [5, 16], we use a 3D-GCN-based [31] graph convolution backbone network to fully exploit the local features in the scene.
@@
Transformer module. Inspired by previous works [6, 24], we use transformers to extract the global context within the feature graph, to boost place recognition performance. The architecture of our devised transformer module is depicted in Fig. 5. To enable the transformer to capture the spatial correlations embedded in the feature graph, we use a feed-forward network to encode the coordinates of the feature graph $p_{\mathrm{feat}}^{i}$ into learnable positional embeddings. We add the positional embeddings to the features and use stacked 3D graph convolution layers for feature fusion. Then we feed the position-encoded features into multi-head attention to fully extract the global spatio-temporal information in the scene. The self-attention mechanism can be formulated as:
@@
TABLE I: Comparision of place recognition performance on the BS, SON, and SQ splits
@@
AnyLoc [3] $\times$ V 80.79 89.76 94.11 97.47 98.74 100.00 90.55 92.07 93.29
OT [6] $\times$ L 67.60 82.75 86.96 92.68 97.22 98.23 96.95 99.39 99.39
@@
SeqOT [18] $\checkmark$ L 78.12 88.78 92.01 97.47 98.48 98.99 98.78 99.39 99.39